what is the magnitude of a point charge due to which the electric feild 30 cm away has the magnitude 2 newton/coulomb

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi,

electrric field due to apoint charge at a distance x from it

                =  1/4¶ε q/ x^2

              2  = 8.9 *10^9  *q  /  (.30)^2 

               q = 2* .09 / 8.9   * 10^-9

                  =  2.02 * 10^-11   C